a. Using the provided information below,

x=33.2, x2=131.67, y=30.78 , y2=116.52 , xy=119.8 , and n=10

i. Find the sample correlation coefficient r







ii. Find the co-efficient of determination


1
Expert's answer
2022-03-17T06:12:33-0400

i.

"r=\frac{n∑xy-∑x∑y}{\sqrt{[n∑x^2-(∑x)^2][[n∑y^2-(∑y)^2]}}"

"=\frac{10(119.8)-33.2(30.78)}{\sqrt{[10(131.667)-33.2^2][10(116.52)-30.78^2}}"

"=\frac{1198-1024.884}{\sqrt{(214.46)(217.7916)}}"

"=\frac{173.116}{\sqrt{46707.58654}}"

"=\frac{173.116}{216.1194}"

"=0.801"


ii. "r^2=0.6416"